From SOWNWiki
Jump to: navigation, search

Meeting (to be) held on 2008/11/20 at 19:00 in Mary Campbell Room

Previous meeting: 13 November 2008 19:00:00 Next meeting: 27 November 2008 19:00:00


  • WiMAX
    • Decided to tell buildings and estates, Daveruss sent draft email to Tim and cc-ed to the committee.
  • Mobile IPv6
    • No progress.
  • 802.1x
    • There may be a trust issue with eduroam traffic once sent from node to ADSL router.
    • No progress.
  • Home node settings
    • No progress.
  • Zepler Roof Trip
    • Removing 802.11b node
      • Sucess! 2.4GHz aerial and the node were removed, the pole was left on the roof.
    • Survey for WiMAX deployment
      • We would like to have a look at the roof of mountbatten.
  • Routing check via Virtual Interaces
    • Interface is set up, but nagios need to be configured to perform pings over it.
  • LDAP Authentication
    • To Migrate or not to migrate, that is the question
[15:37] crwilliams: right - Leth seems to have managed to get LDAP authentication on sown-dev working nicely now
[15:38] crwilliams: the question now is do we update our LDAP entries to reference our old uids
[15:38] crwilliams: or do we update our uids to match our ldap ones?
      • Ok for login and sudo on Dev.
      • For other servers we would require fine granularity controls for them. Groups per-server rather than per permissions level.
      • Auth should definitely be a manual change, not LDAP based.
    • Config Check - Does anyone have more 'pam' experience and can check over /etc/pam.d/common-auth on dev?
      • No-one at meeting, try next week.
    • Any other services ought to use LDAP?
      • Wiki
        • Ask ECS.
  • Todo List
    • Review of tasks
      • Left for Saturday
    • Development of tool
      • Left for Saturday
  • DNS Records
    • (inside ECS)
;; AUTHORITY SECTION:            60562   IN      NS            60562   IN      NS            60562   IN      NS            60562   IN      NS
    • (inside ECS)
;; AUTHORITY SECTION:            60473   IN      NS            60473   IN      NS            60473   IN      NS            60473   IN      NS
    • We need to sort out changing the authority for both of these to ECS permanently
  • Workshop
    • Sort out a consistent name for these sessions
      • Down the pub
    • Tasks
      • Test stuff removed from zepler, as per how it would be deployed.
      • Todo list for people without DPA looking for coding.
      • Contents of the todo list
      • Home Node settings.
  • Services for SOWN
    • SOWN VoIP
      • Becky will have something next week, we are told.
  • SOWN Polo Shirts Order
    • No Progress this week.
  • AOB
    • Equiptment acquired
      • Four new 2.4GHz omni aerials acquired (3x 5dbi, 1x 9dbi).
      • 3x N-Male to N-Male .5m cables.
      • Various Mounting brackets.
    • SJH has updated & upgraded the packages all sown servers.
... more about "2008/11/20"
19:00:00, 20 November 2008 +
20:00:00, 20 November 2008 +